# Histopathology Foundation Models Enable Accurate Ovarian Cancer Subtype Classification

### v1

@@ Methods
The first computational step in the classification pipeline (Figure 3) was tissue segmentation. A significant proportion of most WSIs is non-tissue background which can be discarded using saturation thresholding, where only the pixels with saturation higher than the threshold are labelled as tissue. Otsu thresholding [44] automatically determines the threshold for each image by minimising the variance within the high-saturation and low-saturation groups. Saturation thresholding is computationally efficient, but risks including artifacts such as bubbles, pen marks, and coverslip edges in the foreground. While more robust (and complex) tissue segmentation techniques exist [45, 46], we focused on simple approaches as the attention mechanism in the classification models should learn to ignore any remaining artifacts. We compared the CLAM [24] default static threshold (8/255) to Otsu thresholding with parameters manually adjusted to qualitatively improve the segmentation. The Otsu procedure was found to remove some unstained tissue and artifacts, while also missing some small areas of stained tissue which may have contained diagnostically relevant information (Figure 1).

### v2

@@ Appendix D Augmentation and Normalisation Analysis
Otsu thresholding [58] is applied during tissue detection to automatically determine the saturation threshold for each image by minimising the variance within the separated high-saturation and low-saturation groups. Saturation thresholding is a computationally efficient tissue segmentation approach, but risks including artifacts such as bubbles, pen marks, and coverslip edges in the foreground. While more robust (and complex) tissue segmentation techniques exist [63, 64], we focused on simple approaches as the attention mechanism in the classification models should learn to ignore any remaining artifacts. We compared the CLAM [28] default static threshold (8/255) to Otsu thresholding with parameters manually adjusted to qualitatively improve the segmentation.
Augmentation techniques conversely increase the variability of the training data to allow the model to learn a more general domain. For such large images, training models end-to-end to allow for online data augmentation (adjustments during training) is extremely computationally intensive [66]. Some researchers have attempted to apply online augmentations in the embedding space using generative models [67, 30], though this adds an extra layer of complexity to an already resource-intensive model pipeline. Instead, offline augmentation creates a finite set of augmented versions of the original data, artificially increasing the diversity of training data to a lesser extent than online augmentation. We investigated colour augmentations which adjusted the brightness, contrast, saturation and hue of each patch using parameters from a previous study [68], which we found to create plausibly altered colours (Figure 9).
Extended Results and Discussion
Preprocessing Approach Balanced Accuracy AUROC F1 Score
Baseline 57.1% (53.8-60.4%) 0.893 (0.879-0.907) 0.596 (0.561-0.630)
Reinhard Normalisation 51.3% (48.2-54.4%) 0.872 (0.856-0.887) 0.520 (0.488-0.553)
Macenko Normalisation 57.8% (54.5-61.2%) 0.882 (0.867-0.896) 0.601 (0.567-0.635)
Otsu Thresholding 53.9% (50.6-57.2%) 0.888 (0.873-0.903) 0.566 (0.532-0.600)
Otsu + Macenko 58.0% (54.6-61.4%) 0.882 (0.865-0.898) 0.605 (0.571-0.638)
5x Colour Augmentation 57.4% (54.0-60.7%) 0.888 (0.873-0.902) 0.592 (0.560-0.625)
10x Colour Augmentation 59.1% (55.7-62.4%) 0.891 (0.877-0.905) 0.615 (0.581-0.649)
20x Colour Augmentation 59.1% (55.7-62.4%) 0.892 (0.877-0.905) 0.596 (0.564-0.627)
Table 14: Results of five-fold cross-validation for the ImageNet-pretrained ResNet50 with varied preprocessing approaches. Results are reported as the mean and 95% confidence intervals (in brackets) from 10,000 iterations of bootstrapping. The greatest results are shown in bold.
Preprocessing Approach Balanced Accuracy AUROC F1 Score
Baseline 66.0% (58.1-73.7%) 0.916 (0.873-0.953) 0.634 (0.537-0.726)
Reinhard Normalisation 65.0% (56.6-73.2%) 0.923 (0.881-0.961) 0.632 (0.534-0.727)
Macenko Normalisation 63.0% (54.4-71.5%) 0.915 (0.873-0.951) 0.620 (0.521-0.715)
Otsu Thresholding 65.0% (56.7-73.4%) 0.916 (0.872-0.955) 0.637 (0.542-0.732)
Otsu + Macenko 59.0% (50.3-67.6%) 0.918 (0.878-0.952) 0.577 (0.475-0.674)
5x Colour Augmentation 65.0% (57.0-72.9%) 0.916 (0.876-0.951) 0.630 (0.536-0.725)
10x Colour Augmentation 64.0% (55.9-72.1%) 0.906 (0.864-0.944) 0.616 (0.522-0.710)
20x Colour Augmentation 68.0% (59.7-76.0%) 0.904 (0.861-0.942) 0.657 (0.563-0.750)
Table 15: Results of hold-out testing for the ImageNet-pretrained ResNet50 with varied preprocessing approaches, with predictions generated by an ensemble of the five-fold classification models. Results are reported as the mean and 95% confidence intervals (in brackets) from 10,000 iterations of bootstrapping. The greatest results are shown in bold.
Preprocessing Approach Balanced Accuracy AUROC F1 Score
Baseline 69.2% (58.7-79.7%) 0.956 (0.928-0.980) 0.696 (0.582-0.807)
Reinhard Normalisation 75.8% (65.1-86.0%) 0.968 (0.943-0.986) 0.761 (0.647-0.861)
Macenko Normalisation 74.5% (64.3-84.3%) 0.959 (0.933-0.980) 0.756 (0.648-0.857)
Otsu Thresholding 77.2% (66.4-87.6%) 0.963 (0.937-0.985) 0.797 (0.685-0.895)
Otsu + Macenko 80.5% (70.4-89.9%) 0.983 (0.967-0.995) 0.834 (0.730-0.921)
5x Colour Augmentation 74.9% (63.8-85.6%) 0.966 (0.941-0.986) 0.762 (0.647-0.866)
10x Colour Augmentation 76.1% (65.0-86.6%) 0.962 (0.935-0.983) 0.768 (0.659-0.869)
20x Colour Augmentation 80.0% (69.2-90.0%) 0.973 (0.953-0.989) 0.806 (0.706-0.897)
Table 16: Results of external validation on the Transcanadian Study dataset for the ImageNet-pretrained ResNet50 with varied preprocessing approaches, with predictions generated by an ensemble of the five-fold classification models. Results are reported as the mean and 95% confidence intervals (in brackets) from 10,000 iterations of bootstrapping. The greatest results are shown in bold.
Preprocessing Approach Balanced Accuracy AUROC F1 Score
Baseline 52.4% (49.5-55.1%) 0.868 (0.847-0.889) 0.412 (0.380-0.444)
Reinhard Normalisation 51.0% (47.7-54.3%) 0.870 (0.850-0.888) 0.392 (0.350-0.437)
Macenko Normalisation 45.9% (41.8-50.0%) 0.837 (0.814-0.860) 0.407 (0.360-0.455)
Otsu Thresholding 54.7% (51.9-57.6%) 0.883 (0.864-0.901) 0.440 (0.401-0.482)
Otsu + Macenko 44.4% (40.7-48.3%) 0.840 (0.816-0.862) 0.388 (0.347-0.432)
5x Colour Augmentation 51.7% (48.4-54.8%) 0.867 (0.845-0.887) 0.401 (0.363-0.441)
10x Colour Augmentation 51.1% (47.8-54.2%) 0.877 (0.856-0.897) 0.404 (0.367-0.443)
20x Colour Augmentation 51.4% (48.4-54.4%) 0.874 (0.853-0.893) 0.391 (0.352-0.433)
Table 17: Results of external validation on the OCEAN Challenge dataset for the ImageNet-pretrained ResNet50 with varied preprocessing approaches, with predictions generated by an ensemble of the five-fold classification models. Results are reported as the mean and 95% confidence intervals (in brackets) from 10,000 iterations of bootstrapping. The greatest results are shown in bold.
The Otsu procedure was found to remove some unstained tissue and artifacts, while also missing some small areas of stained tissue which may have contained diagnostically relevant information (Figure 11). It was unclear from the results whether this was beneficial or harmful overall, with Otsu thresholding giving inconsistent performance when compared to the baseline.